OpenStudy (anonymous):

The equation of the circle with center (3, -2) and radius 7 is

OpenStudy (lgbasallote):

equation of a circle is \[(x-h)^2 + (y-k)^2 = r^2\] where h is the x coordinate of the center k is the y coordinate of the center r is radius

OpenStudy (anonymous):

is the answer x + y = 51

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (lgbasallote):

maybe a demonstration will help...

OpenStudy (lgbasallote):

EXAMPLE: center is at (2,3) and radius is 4 \[(x-h)^2 + (y-k)^2 = r^2\] \[(x - 2)^2 + (y-3)^2 = 4^2\] \[(x-2)^2 + (y-3)^2 = 16\] this is now the equation for the one i gave

OpenStudy (anonymous):

(x - h)2 +(y-k)2 = r2 (x-3)2 + (y- -2)2 = 49 x-9 + y + 4 = 49 x-5 + y = 49 x+y = 54

OpenStudy (anonymous):

(x-3)2 + (y- -2)2 = 49 is this supposed to be the answer

OpenStudy (lgbasallote):

simplify it further... what is y-(-2)

OpenStudy (anonymous):

y+4

OpenStudy (lgbasallote):

there...that's the answer \[(x-3)^2 + (y+4)^2 = 49\]